OpenStudy (anonymous):

solve and round to the nearest hundreth 3x^2+5x+7=0

jhonyy9 (jhonyy9):

there are complex roots

OpenStudy (anonymous):

i know I dont get it

OpenStudy (anonymous):

dont you gotta combine like terms first?

OpenStudy (anonymous):

\[3x^2+5x+7=0 \\ b^2 - 4ac =25- 4(3)(7)=-5 \] Th roots are complex

OpenStudy (anonymous):

\[\{x=-0.8-1.3 i,x=-0.8+1.3 i\} \]